### v1

# CoAR: Concept Injection into Autoregressive Models for Personalized Text-to-Image Generation

Abstract
The unified autoregressive (AR) model excels at multimodal understanding and generation, but its potential for customized image generation remains underexplored. Existing customized generation methods rely on full fine-tuning or adapters, making them costly and prone to overfitting or catastrophic forgetting. In this paper, we propose CoAR, a novel framework for injecting subject concepts into the unified AR models while keeping all pre-trained parameters completely frozen. CoAR learns effective, specific subject representations with only a minimal number of parameters using a Layerwise Multimodal Context Learning strategy. To address overfitting and language drift, we further introduce regularization that preserves the pre-trained distribution and anchors context tokens to improve subject fidelity and re-contextualization. Additionally, CoAR supports training-free subject customization in a user-provided style. Experiments demonstrate that CoAR achieves superior performance on both subject-driven personalization and style personalization, while delivering significant gains in computational and memory efficiency. Notably, CoAR tunes less than 0.05% of the parameters while achieving competitive performance compared to recent Proxy-Tuning. Code: https://github.com/KZF-kzf/CoAR.
Figure 1: Visualization of samples generated by our proposed CoAR, showcasing subject personalization (left), where a few reference images are provided, and style personalization (right), where the subject is rendered in user-specified artistic styles without additional training.
Introduction
Subject-driven image generation (Chen et al. 2023b; Ruiz et al. 2023; Li, Li, and Hoi 2023; Shin et al. 2025) aims to generate images of a specific subject or concept in new scenes, given only a few reference images. This capability has found applications in virtual try-on (Ye et al. 2025; Zhang et al. 2024), personalized content creation (Liu et al. 2024b), and advertising design (Chen et al. 2023b).
Most of the existing approaches, such as DreamBooth (Ruiz et al. 2023) and its extensions (Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023), customize generation by adapting the distribution of pre-trained diffusion models to match the subject distribution through full fine-tuning or parameter-efficient techniques like LoRA (Hu et al. 2022). These methods typically associate the subject with a unique identifier [V], enabling personalized generation via prompt injection during inference. However, such techniques are predominantly designed for diffusion models (Esser et al. 2024; Rombach et al. 2022; Podell et al. 2023) and are not directly transferable to multimodal autoregressive (AR) models. Applying these methods naively to AR models poses two challenges.
  • Risk of overfitting. AR models are pretrained on large-scale multimodal corpora and expected to support both understanding and generation. In contrast, the target subject typically has only a few reference images (3–5). Such scarce data makes it extremely difficult to adapt the pre-trained models to the subject distribution, resulting in a risk of overfitting and catastrophic forgetting of the pre-trained knowledge.
  • Inefficiency. Extensively adjusting the pretrained model parameters to learn a single concept with the limited number of images per subject is intuitively inefficient.
We argue that instead of fine-tuning the pretrained model on limited data to shift its distribution toward a target subject, it is more effective to inject subject concepts directly into the model’s existing distribution, thereby fully leveraging the generative capabilities of large models. To this end, we introduce Concept Injection for AR Models (CoAR), an extremely efficient framework for customized generation in multimodal AR models. CoAR adopts a Layer-wise Multimodal Context Learning (LMCL) strategy, encoding subject-specific information through a small number of learnable prompts without modifying any pretrained parameters. To guide learning while preserving the model’s original capabilities, we propose a Dual Prior Preservation (DPP) loss that constrains training. Additionally, we design a Context-Aware Self-Regularization (CASR) loss to improve re-contextualization stability and output quality. Built upon this foundation, CoAR also supports training-free customization, enabling arbitrary subject–style combinations during inference while maintaining strong re-contextualization performance. With a frozen backbone and only minimal parameter tuning, CoAR effectively injects the specific subject concept into the AR model and mitigates overfitting through multi-faceted regularization.
Extensive experiments demonstrate the superiority of CoAR in subject fidelity, instruction alignment, and parameter efficiency. Compared to naive DreamBooth adaptation in AR models, CoAR achieves significant performance gains while training approximately fewer than 0.1M parameters. Against existing AR-based customization methods, CoAR achieves competitive or superior results while training with no more than 0.15% of the parameters used by other approaches. For stylization tasks, our training-free method achieves subject identity preservation and high-fidelity stylization across diverse scenes, matching the modern LoRA-based competitor (Shah et al. 2024) that requires additional training. Our main contributions are summarized as follows:
  • We propose a novel framework called CoAR for AR customization. Unlike mainstream LoRA-based approaches, CoAR injects subject concepts into minimal context tokens without modifying pre-trained parameters.
  • We propose a training-free customization strategy that enables arbitrary subject–style combinations during inference. By concatenating subject and style context tokens, our method achieves high-quality stylization while preserving subject identity and contextual consistency without any additional training. Some generated samples are shown in Fig.˜1
  • Experimental results demonstrate that CoAR achieves competitive or superior performance on subject-driven personalized generation, while requiring substantially fewer trainable parameters. Furthermore, it achieves performance comparable to modern diffusion-based approaches in subject-style personalized generation.
Related Works
Autoregressive Image Generation
In recent years, autoregressive (AR) image generation has rapidly evolved from pixel-level sequential modeling to token-level generation and further to multi-modal unified architectures. Early approaches (Van den Oord et al. 2016; Van Den Oord, Kalchbrenner, and Kavukcuoglu 2016) operated directly in the pixel space, which imposed heavy computational demands for generating high-resolution images. The VQ-VAE family (Ramesh et al. 2021; Van Den Oord, Vinyals et al. 2017) subsequently discretizes images into tokens, followed by training a Transformer (Vaswani et al. 2017) model using the Next Token Prediction (NTP) task to generate images. Recent research has explored two major directions: One is Visual autoregressive (VAR) models (Tian et al. 2024; Han et al. 2025). VAR models adopt a next-scale prediction framework that significantly improves image quality. The other is multi-modal unified AR models (Team 2024; Liu et al. 2024a; Xin et al. 2025). Chameleon (Team 2024) employs an early-fusion strategy to mix text and image tokens into a single sequence, enabling unified text-to-image, captioning, and inpainting generation. Lumina-mGPT (Liu et al. 2024a) further incorporates progressive FP-SFT, substantially enhancing high-resolution image synthesis. Lumina-mGPT 2.0 (Xin et al. 2025) approaches state-of-the-art diffusion models within a pure AR framework while maintaining strong multi-task generalization. These unified multi-modal AR models demonstrate strong potential as a foundation for versatile and scalable vision-language generation systems.
Figure 2: Overview of the proposed CoAR framework for subject-driven personalization in multi-modal autoregressive models. (a) Layerwise Multimodal Context Learning, where learnable context tokens are injected into multiple Transformer layers for concept representation. (b) Dual Prior Preservation (DPP) regularizes the customized distribution against the pre-trained model to mitigate overfitting and language drift. (c) Context-Aware Self-Regularization (CASR) initializes and constrains context tokens towards the subject embedding space to enhance fidelity and re-contextualization. (d) Training-free subject–style composition by directly combining subject and style tokens to enable flexible customized generation.
Subject-Driven Image Generation
Subject-driven image generation aims to capture the concept of a specific subject from a limited set of reference images and generalize it to novel contexts. Existing approaches can broadly be categorized into diffusion-based and AR-based methods. Diffusion-based approaches (Ruiz et al. 2023; Chen et al. 2023a; Han et al. 2023; Kumari et al. 2023; Voynov et al. 2023; Gal et al. 2022) typically adapt the distribution of a pre-trained model to that of the target subject via either full fine-tuning or low-rank adaptation (LoRA) (Hu et al. 2022). During inference, a special identifier is employed to sample from the adapted distribution, allowing for customized generation. As the potential of AR models in image generation continues to be exploited, recent studies (Chung et al. 2025; Zheng et al. 2025; Sun et al. 2025; Wu et al. 2025) have begun to investigate subject-driven customization specifically tailored for AR models. For VAR-based models, ARBooth (Chung et al. 2025) introduces selective layer tuning to reduce complexity and employs prior distillation to mitigate the issue of language drift. FreeLoRA (Zheng et al. 2025) proposes a training-free multi-subject customization approach for VAR models. For multimodal AR models, PersonalAR (Sun et al. 2025) validates the combination of DreamBooth (Ruiz et al. 2023) and TI (Gal et al. 2022) within a LoRA framework on Lumina-mGPT. Proxy-Tuning (Wu et al. 2025) further enhances capability in subject-driven image generation by leveraging diffusion models as proxies.
These approaches follow the diffusion-based customization paradigm, adapting the generation backbone to specific subject distributions. A concurrent work, Yo’Chameleon (Nguyen et al. 2025), also introduces context optimization for multimodal generation of the auto-regressive model but suffers from overfitting and language drift, leading to reduced quality and subject fidelity. In contrast, CoAR represents subject or style information using a compact set of learnable context tokens. This eliminates the need to update backbone parameters while preserving the model’s generative capacity. By decoupling customization from backbone adaptation, CoAR offers an effective alternative to prior approaches.
Methodology
In this work, we propose CoAR, a simple yet effective framework for subject-driven customization in multi-modal autoregressive (AR) models. The overall CoAR framework is shown in Fig. 2, consisting of:
  • Layer-wise Multi-modal Contextual Learning: injecting learnable multimodal tokens into multiple Transformer layers.
  • Dual Prior Preservation: regularizing training with the frozen pre-trained model.
  • Context-Aware Self-Regularization: initializing and constraining context tokens with subject embeddings to prevent overfitting.
Preliminary
We first introduce the AR: in multimodal autoregressive (AR) text-to-image generation, both text and images are represented as sequences of discrete tokens drawn from a shared vocabulary $\mathcal{V}$. A text prompt is encoded as:
$\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L}\},\quad\mathbf{y}_{l}\in\mathcal{V},$ (1)
and the target image is represented as
$\mathbf{x}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},\quad\mathbf{x}_{t}\in\mathcal{V},$ (2)
where $L$ and $T$ denote the lengths of the text and image token sequences, respectively.
AR model generates the image autoregressively by iteratively sampling the next token conditioned on the text and all previously generated tokens:
$\mathbf{x}_{t}\sim p_{\theta}(\mathbf{x}_{t}\mid\mathbf{x}_{<t},\mathbf{y}),\quad t=1,\dots,T,$ (3)
where $\mathbf{x}_{<t}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{t-1}\}$ and $\theta$ are the model parameters.
Training follows the next-token prediction paradigm, minimizing the negative log-likelihood of the ground-truth token sequence conditioned on the text prompt:
$\mathcal{L}_{\text{NTP}}=-\sum_{t=1}^{T}\log p_{\theta}(\mathbf{x}_{t}\mid\mathbf{x}_{<t},\mathbf{y}),$ (4)
where $\mathbf{x}_{t}$ denotes the token at time step $t$, $\mathbf{x}_{<t}$ represents the preceding tokens, and $y$ is the input text prompt. This objective is implemented using the standard cross-entropy loss over the shared vocabulary $\mathcal{V}$.
Layerwise Multimodal Context Learning Strategy
We introduce a Layerwise Multimodal Contextual Learning (LMCL) strategy to capture the target subject concept while keeping the pre-trained model entirely frozen, requiring only a small number of learnable parameters for effective customization.
Specifically, for all images from the same subject, we define a shared set of multimodal learnable tokens as the representation of the subject concept:
$\mathbf{P}=\{\mathbf{p}_{{[v]}},\mathbf{p}_{{I}}\},$ (5)
where $\mathbf{p}_{{[v]}}\in\mathbb{R}^{1\times N\times D}$ denotes the learnable tokens for the text modality with the number of insertion layers $N$ and the token embedding dimension $D$. Similarly, $\mathbf{p}_{{I}}$ represents the learnable tokens for the image modality with the same shape $1\times N\times D$.
For each subject image, we construct a textual template using a shared prompt a photo of [V] [Class], where [V] is a placeholder representing the subject identity and [Class] is the class name. Given the original token sequence at the $i_{th}$ Transformer layer $\mathbf{U}_{i}$ (omitting special start/end tokens for simplicity):
$\mathbf{U}_{i}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{y}_{L},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\}.$ (6)
As shown in Fig.˜2(a), we insert the learnable tokens as:
$\mathbf{U}_{i}^{\prime}=\{\mathbf{y}_{1},\mathbf{y}_{2},\dots,\mathbf{p}_{{[v]}}^{(i)},\dots,\mathbf{y}_{L},\mathbf{p}_{{I}}^{(i)},\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}\},$ (7)
where $\mathbf{p}_{{[v]}}^{(i)}$ replaces the placeholder [V] in the text sequence, and $\mathbf{p}_{{I}}^{(i)}$ is inserted before the first image token.
Finally, we optimize the representation of these multimodal tokens using Eq. (4) to progressively inject the concept of the specific subject into them. By learning only $\{\mathbf{p}_{[v]},\mathbf{p}_{I}\}$ while keeping the backbone frozen, LMCL achieves efficient subject-driven customization with minimal parameter overhead.
Dual Prior Preservation
To mitigate language drift and encourage generation diversity, we propose Dual Prior Preservation (DPP) to make the prior preservation introduced in DreamBooth (Ruiz et al. 2023) compatible with training of the multimodal AR model. We first generate a small set (6–8) of class images using the pre-trained model with the prompt “a photo of a [Class]”. These generated images, along with their corresponding text prompts, are tokenized into an input sequence $\mathbf{U}_{{cls}}$. We then perform a dual forward pass on $\mathbf{U}_{{cls}}$:
  • Forwarding the original $\mathbf{U}_{{cls}}$ through the pretrained model to obtain a zero-shot output distribution ${\text{logits}}_{{zs}}$.
  • Inserting layerwise image context tokens into $\mathbf{U}_{{cls}}$ yields $\mathbf{U}_{{cls}}^{\prime}$ and forward through the current model to obtain $\text{logits}_{{prior}}$.
The DPP loss is computed as:
$\mathcal{L}_{{DPP}}$ $=\lambda_{1}\cdot\mathcal{L}_{{NTP_{cls}}}({\text{logits}}_{{prior}},\text{labels}_{cls})$ (8)
$+\lambda_{2}\cdot D_{{KL}}({\rm{logits}}_{{zs}}\parallel{\rm{logits}}_{{prior}}),$
where $\mathcal{L}_{{NTP_{cls}}}$ denotes the Next Token Prediction (NTP) loss on the class images in Eq. (4), $\text{labels}_{cls}$ are the ground-truth visual tokens of the class images, $D_{{KL}}$ is the Kullback–Leibler (KL) divergence measuring the distance between the two output distributions, which is formulated as:
$\begin{aligned} D_{{KL}}({\rm{logits}}_{{zs}}\parallel{\rm{logits}}_{{prior}})&=\\ \sum_{v\in\mathcal{V}}{\rm{logits}}_{{zs}}(v)\log&\frac{{\rm{logits}}_{{zs}}(v)}{{\rm{logits}}_{{prior}}(v)}\end{aligned},$ (9)
and $\lambda_{1},\lambda_{2}$ are coefficients balancing the NTP and KL regularization terms.
The dual-path design extends DreamBooth’s prior preservation from continuous diffusion latents to discrete token modeling, making it compatible with multi-modal AR frameworks.
Context-Aware Self-Regularization
To address overfitting and improve re-contextualization in subject-driven customization, we propose the Context-Aware Self-Regularization (CASR). As shown in Fig.˜2(c), given the reference images of the specific subject, we feed all samples into the frozen pre-trained model and compute the average embedding of all image tokens in each layer $i$, denoted as $\mathbf{E}_{{subject}}^{(i)}$. We initialize the learnable image context token $\mathbf{p}_{I}^{(i)}$ with $\mathbf{E}_{{subject}}^{(i)}$, enabling training to start from a representation closer to the subject’s distribution and accelerating convergence. The CASR loss is defined as:
$\mathcal{L}_{{CASR}}=\frac{1}{N}\sum_{i=1}^{N}\big{\|}\mathbf{p}_{I}^{(i)}-\mathbf{E}_{{subject}}^{(i)}\big{\|}^{2},$ (10)
where $\|\cdot\|^{2}$ denotes the squared $\ell_{2}$-norm measuring the Euclidean distance and $N$ is the total number of insertion layers.
CASR regularizes the context tokens by constraining them towards the subject’s embedding space, effectively preventing overfitting and significantly improving re-contextualization performance.
Overall Objective
The final training objective combines the autoregressive next-token prediction with the proposed regularization terms:
$\mathcal{L}_{obj}=\mathcal{L}_{{NTP}}+\alpha\mathcal{L}_{{DPP}}+\beta\mathcal{L}_{{CASR}},$ (11)
where $\mathcal{L}_{{NTP}}$ denotes the next-token prediction loss on the subject images, $\alpha$ and $\beta$ are two factors that balance the contribution of the DPP and CASR losses, respectively.
Method DINO CLIP-I CLIP-T Trainable Params Training Stage
Real Imges 0.774 0.885 N/A N/A N/A
TI (Ruiz et al. 2023) 0.569 0.780 0.255 N/A
DreamBooth(SD) (Ruiz et al. 2023) 0.668 0.803 0.305 N/A
DreamBooth (Imagen) (Ruiz et al. 2023) 0.696 0.812 0.306 N/A One-Stage
DreamBooth (Lumina-mGPT) (Wu et al. 2025) 0.5088 0.6752 0.2956 N/A
DreamBooth (Lumina-mGPT)+LoRA (Wu et al. 2025) 0.5338 0.6974 0.2652 N/A
PersonalAR (Lumina-mGPT)+LoRA (Sun et al. 2025) 0.657 0.778 0.312 50.4M
PersonalAR (Lumina-mGPT) (Sun et al. 2025) 0.671 0.805 0.302 1610.6M Two-Stage
Proxy-Tuning (Lumina-mGPT/SD3.5)+LoRA (Wu et al. 2025) 0.7523 0.8091 0.3123 142.6M
CoAR (Lumina-mGPT) (Ours) 0.7226 0.8151 0.3184 0.073M One-Stage
Table 1: Performance comparison of our CoAR and the recent competitors on the DreamBench dataset (Ruiz et al. 2023). The best results are marked in bold.
Figure 3: Qualitative comparison of subject-driven personalization on the DreamBench benchmark, selecting PersonalAR and Proxy-Tuning for comparison. The results for both competitors are taken directly from their original publications.
Customize Any Subject in Any Style
Since CoAR leaves the pre-trained AR model parameters untouched, it fully leverages the strong multimodal understanding and generative capacity of autoregressive models, which naturally support zero-shot subject–style composition. This design enables training-free customization of arbitrary subjects in arbitrary styles simply by concatenating their respective context tokens, allowing for flexible, efficient, and scalable rendering without fine-tuning. To further ensure clean and controllable generation, we introduce an Identity Mask during inference, as shown in Fig.˜2(d), which explicitly restricts attention flow between subject and style tokens. This mechanism prevents semantic interference and ensures that each token type contributes distinctly and independently to the final generation.
Experiments
Experimental Details
Figure 4: Re-contextualization with our CoAR on DreamBench (Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
Benchmark. We evaluate subject-driven customization using the DreamBench dataset (Ruiz et al. 2023), which contains 30 subjects (21 objects, 9 animals), each with 3–5 reference images and 25 test prompts. Objects are tested with 20 recontextualization and 5 property modification prompts, while animals use 10 recontextualization, 10 accessorization, and 5 property modification prompts. For stylization, we adopt the StyleDrop dataset (Sohn et al. 2023) and follow ZipLoRA (Shah et al. 2024), training with a single reference image per style. Additionally, we evaluate subject-style hybrid generation by combining DreamBench subjects with StyleDrop styles.
Implementation Details. All experiments are run on a single H800 GPU using Lumina-mGPT-7B FP-SFT (Liu et al. 2024a) as the backbone. For subject-driven personalization, we insert one text and one image token per layer into the first 9 layers. For stylization with a single reference image, the same token configuration is applied to the first 3 layers. The model is trained with a learning rate of 1e-2 and batch size 1, keeping all other Lumina-mGPT hyperparameters unchanged. We set $\alpha$ = 1e-2, $\beta$ = 5e-4, and both $\lambda_{1}$ and $\lambda_{2}$ to 0.5. Each subject and style is trained for 1000 and 600 steps, respectively.
CoAR (Ours) B-LoRA ZipLoRA
Subject-alignment 0.604 0.579 0.655
Style-alignment 0.594 0.505 0.597
Text-alignment 0.308 0.258 0.272
Table 2: Comparison results of our CoAR and two competitors on the style personalization task.
Evaluation Metrics. For subject-driven customization tasks, we follow established protocols and employ CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) for subject fidelity evaluation, along with CLIP-T (Radford et al. 2021) for prompt fidelity assessment. Specifically, CLIP-I computes the average cosine similarity between CLIP embeddings of real and generated images. DINO measures the average cosine similarity of ViT-S/16 DINO embeddings between real and generated images. CLIP-T evaluates the cosine similarity between CLIP text embeddings of prompts and CLIP image embeddings of generated outputs. For stylization tasks, we adopt metrics from ZipLoRA (Shah et al. 2024), which use CLIP features for both style and text alignment, and DINO features for subject alignment.
Competitors. For the subject-driven personalization, we compare our approach with several baselines: Textual Inversion (TI) (Gal et al. 2022), DreamBooth (Ruiz et al. 2023), and two DreamBooth-based competitors (Wu et al. 2025), which share the same autoregressive backbone as our method. We also include two recent approaches specifically designed for multimodal autoregressive models (Sun et al. 2025; Wu et al. 2025). For the style personalization, we compare with recent diffusion-based methods: ZipLoRA (Shah et al. 2024) and B-LoRA (Frenkel et al. 2024).
Main Results
Quantitative Results of Subject-Driven Personalization.
Table 1 shows the comparison results of our CoAR and the competitors on the DreamBench dataset. From the results, we have the following observations. First, CoAR achieves superior performance across multiple dimensions. In the CLIP-I metric, CoAR achieves a leading score of 0.8151, surpassing the second-best DreamBooth (Imagen) by 0.4%, demonstrating its exceptional image fidelity. For CLIP-T, CoAR sets a new benchmark at 0.3184, outperforming the second-best method, Proxy-Tuning by 0.29%, highlighting its superior cross-modal alignment capability. While CoAR ranks second on the DINO metric, it is substantially more efficient and user-friendly than the top-performing method, Proxy-Tuning, which requires training an additional diffusion model and generating hundreds to a thousand images per subject for data augmentation.
Quantitative Results of Style Personalization.
Table 2 presents a comparison between our CoAR model and two baselines: B-LoRA (Frenkel et al. 2024) and ZipLoRA (SDXL) (Shah et al. 2024). The results demonstrate that CoAR outperforms B-LoRA across all three evaluation metrics, achieving the highest text-alignment score among all competitors. While CoAR trails ZipLoRA slightly in subject alignment (0.604 vs. 0.655), it achieves comparable style preservation (0.594 vs. 0.597), superior text alignment (0.308 vs. 0.272), and does so without requiring any additional training.
Qualitative Results of Subject-Driven Personalization.
Fig.˜3 illustrates some generated samples of our CoAR compared with two most related competitors, i.e., PersonalAR (Sun et al. 2025) and Proxy-Tuning (Wu et al. 2025). From the results, we observe that images generated by CoAR exhibit superior prompt fidelity. For instance, given the instruction “wearing a black top hat and a monocle", neither PersonAR nor Proxy-Tuning successfully generates the “monocle", whereas our CoAR does. Moreover, our generated images better adhere to physical plausibility. In the second row of PersonAR, the generated vase and backpack are both floating, which is unrealistic, while such artifacts are absent in our results. More qualitative results are provided in the supplementary materials.
Figure 5: Qualitative comparison of our CoAR and LoRA-based method for style personalization tasks.
Qualitative Results of Style Personalization.
As shown in Fig.˜5, CoAR demonstrates superior performance compared to B-LoRA. While ZipLoRA preserves more subject details, such as color, CoAR exhibits stronger adherence to instructions when generating in novel scenes (e.g., “in the jungle” or “floating in water”). Additional qualitative results on style customization are presented in Fig.˜4, showing that CoAR effectively preserves the given style when generalizing the reference subject to novel scenes.
Ablation Studies
Impacts of Different Losses. To assess the contribution of each loss component in our framework, we perform an ablation study by incrementally applying the LMCL, DPP, and CASR losses and evaluating performance on DINO, CLIP-I, and CLIP-T metrics, as shown in Tab. 3. Only applying LMCL yields baseline scores of 0.6610 (DINO), 0.7647 (CLIP-I), and 0.3096 (CLIP-T). Introducing DPP leads to consistent improvements across all metrics, increasing DINO to 0.6902 and CLIP-I to 0.7919. Although CLIP-T slightly decreases to 0.2968, the overall alignment benefits are clear. When CASR is further added, we observe the best overall performance: 0.7226 (DINO), 0.8151 (CLIP-I), and 0.3184 (CLIP-T), confirming that each loss component contributes positively to the alignment quality and subject fidelity.
LMCL DPP CASR DINO CLIP-I CLIP-T
0.6610 0.7647 0.3096
0.6902 0.7919 0.2968
0.7226 0.8151 0.3184
Table 3: Impacts of different losses on the subject-driven personalization task.
Impacts of Layer Numbers. To evaluate the impacts of depth of multimodal context token insertion, we insert tokens at layers 1, 3, 9, and 24. As illustrated in Fig.˜6. CLIP-T remains relatively stable across depths, while CLIP-I and DINO follow a rise-then-fall trend. Notably, inserting tokens at deeper layers (e.g., layer 24) leads to a decline in subject fidelity. We attribute this to overfitting caused by limited reference data, which becomes more pronounced as token influence shifts to higher-level semantic layers.
More experimental results are provided in the appendix.
Figure 6: Impacts of insertion depth of multimodal context tokens on performance.
Conclusion
In this paper, we introduced CoAR, a novel and efficient framework for subject-driven and style-aware customization in multimodal autoregressive (AR) models. Unlike conventional fine-tuning or LoRA-based approaches, CoAR injects subject concepts directly into the pretrained distribution through a small number of learnable context tokens, avoiding any modification to pretrained parameters. Extensive experiments across personalization and stylization tasks demonstrate that CoAR achieves superior subject fidelity, instruction alignment, and style preservation while training fewer than 0.1M parameters, significantly more parameter-efficient than existing approaches. Moreover, our training-free strategy matches or surpasses modern LoRA-based diffusion methods, highlighting the potential of multimodal AR models for scalable and flexible personalization.
Appendix
Appendix A Training and Evaluation Details
Training Details. The training details of CoAR on different datasets are shown in Table˜4. For the style personalization task, only the provided single reference image is utilized. Consequently, the hyperparameters for Dual Prior Preservation (DPP) are not applied.
Evaluation Details. For the subject-style generation task, we selected 8 subjects from DreamBench (Ruiz et al. 2023), including 4 animals and 4 objects, as well as 6 styles from StyleDrop (Sohn et al. 2023). These yielded a total of 48 unique subject-style combinations, on which we evaluated our CoAR, ZipLoRA (Shah et al. 2024), and B-LoRA (Frenkel et al. 2024). For each combination, we assigned 10 distinct recontextualization prompts for both objects and living subjects, as illustrated in Table˜5.
DreamBench StyleDrop
lr 1e-2 1e-2
batch size 1 1
resolution 768 768
$\lambda_{1}$ 0.5 N/A
$\lambda_{2}$ 0.5 N/A
$\alpha$ 1e-2 N/A
$\beta$ 5e-4 5e-4
Training Steps 1000 600
Context token layers 9 3
Table 4: Training settings of CoAR on DreamBench(Ruiz et al. 2023) and StyleDrop (Sohn et al. 2023) datasets.
subject context
objects in the snow
on the beach
with a city in the background
on top of a dirt road
on top of green grass
in the jungle
on the mountain
floating in water
on a picnic table
on top of a wooden floor
living subjects wearing a hat
with a crown
riding a bicycle
sleeping
in a boat
in the jungle
on the mountain
floating in water
on a picnic table
playing with a ball
Table 5: Prompt templates categorized by subject type.
Appendix B Details of Parameter Calculation
For PersonalAR (Sun et al. 2025), we directly reference the parameter count reported in its original paper. As for Proxy-Tuning (Wu et al. 2025), since the paper does not provide explicit parameter statistics and the code is not publicly available, we provide a coarse estimate based on the hyperparameter configurations described in the paper.
Proxy-Tuning is a two-stage approach where a diffusion model is first trained, followed by an autoregressive (AR) model that is trained on data generated by the diffusion model. Both models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al. 2022). In our estimation, the diffusion model is based on Stable Diffusion 3.5 (SD3.5 Large) (Esser et al. 2024), while the AR model is Lumina-mGPT 7B (Liu et al. 2024a). We estimate the total number of trainable LoRA parameters based on the typical LoRA application settings.
LoRA Parameter Calculation
For a weight matrix $W\in\mathbb{R}^{d\times d}$, LoRA introduces two trainable matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times d}$, where $r$ is the LoRA rank. The number of trainable parameters for each projection is:
$\text{LoRA params}=2dr$
We assume LoRA is applied to all attention projections: $W_{q}$, $W_{k}$, $W_{v}$, and $W_{o}$.
Diffusion Model (SD3.5)
SD3.5 is based on the DiT (Peebles and Xie 2023) architecture, with 24 Transformer (Vaswani et al. 2017) blocks. Each block includes one self-attention and one cross-attention module. Assuming a hidden size $d=3072$, the LoRA (rank=64) parameters per attention module are:
$4\times 2\times d\times r$ $=8dr$ (12)
[…]
There are a total of $24\times 2=48$ attention modules, leading to:
[…]
AR Model (Lumina-mGPT 7B)
For the AR model, we assume a hidden size $d=4096$ and 32 Transformer layers. Applying LoRA (rank = 64) to the four attention projections in each layer results in:
$4\times 2\times d\times r$ $=$ (14)
$=8dr$
[…]
leading to:
[…]
Total Parameter Count
Summing both components, the total number of trainable LoRA parameters under Proxy-Tuning is approximately:
[…]
This estimate assumes LoRA is applied to all attention projections and excludes feed-forward layers. Including FFNs would increase the total parameter count further, but standard LoRA configurations typically restrict adaptation to attention modules only.
Appendix C Additional Experiments
Ablation
As shown in Fig.˜7, we first analyze the results using only Layer-wise Multi-modal Contextual Learning (LMCL), and then with the addition of the Dual Prior Preservation (DPP) loss. With the DPP constraint, the model better preserves the subject’s identity. However, when recontextualizing the subject in the scene “on a cobblestone street”, it still struggles to accurately render the background. In contrast, incorporating Context-Aware Self-Regularization (CASR) enables the model to successfully generalize to the new scene.
Fig.˜8 provides a visual illustration of the effect of our proposed CASR loss. When the loss weight is too small, the model tends to overfit: although subject fidelity remains high, generalization to novel contexts is limited. Conversely, when the weight is too large, the context tokens become under-optimized, leading to a significant degradation in subject fidelity.
Figure 7: Visualization of the effects of different losses on the generated samples.
Figure 8: Visualization of CASR.
Figure 9: Qualitative results of subject-driven customization with CoAR.
Figure 10: Qualitative results of subject-driven customization with CoAR.
Additional Qualitative Results
Figs.˜9, 10, 12, 11 and 13 presents additional qualitative results of subject-driven personalization and style personalization.
Figs.˜9 and 10 demonstrate that given only a few reference images, CoAR successfully performs subject-driven generation across various tasks, including recontextualization (e.g., novel backgrounds), property modification (e.g., color and shape), and accessorization (e.g., adding glasses or outfits). The results demonstrate strong subject fidelity, adaptability to novel contexts, and fine-grained controllability.
Fig.˜12 presents examples of image generation conditioned on different artistic styles. Each row represents a different subject, while each column corresponds to a specific style. The generated images highlight the model’s ability to retain subject identity while adapting visual attributes such as texture, color palette, and rendering technique to match the target style.
Fig.˜13 illustrates the effects of diverse contextual prompts on image generation for various subjects across different visual styles. Each row corresponds to a distinct subject, while each column varies the context.
Figure 11: CoAR demonstrates effective style personalization conditioned on only a single reference image.
Figure 12: Style-conditioned image generation across subjects without contextual prompts.
Figure 13: Impact of contextual prompts on image generation across subjects and styles.